OpenStudy (anonymous):

X+5 over 3 = x-3 over 4

OpenStudy (anonymous):

i posted this in the other than, does it look like this? \[\frac{ x + 5 }{ 3 } = \frac{ x - 3 }{ 4 }\]

OpenStudy (anonymous):

what can we do with that?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

find x?

OpenStudy (anonymous):

Common denominator rite?

OpenStudy (anonymous):

ahhhhhhh!!

OpenStudy (anonymous):

hahahahaha julie u are truely smart :]

OpenStudy (anonymous):

so the lcd is obviously 12 which gets us?

OpenStudy (anonymous):

Im lost

OpenStudy (anonymous):

ok so the Least common denominator of 3 and 4 is 12 so you have to multiply \[\frac{ x + 5 }{ 3 } * \frac{ 4 }{ 4 }\]

OpenStudy (anonymous):

20x over 12?

OpenStudy (anonymous):

yah. that rite.

OpenStudy (anonymous):

no \[\frac{ 4x + 20}{12}\] but i see where you could have done the 20x

OpenStudy (anonymous):

Ooo wats nxt?

OpenStudy (anonymous):

i mean the solution @gaara438125

OpenStudy (anonymous):

so the other side then becomes\[\frac{ 3x - 9 }{ 12 }\] because we do \[\frac{ x - 3 }{ 4 } * \frac{ 3 }{ 3 }\] so it looks like \[\frac{ 4x + 20 }{ 12 } = \frac{ 3x - 9 }{ 12 }\]

OpenStudy (anonymous):

multiply both sides by 12 to get what?

OpenStudy (anonymous):

Idk hw to do that

OpenStudy (anonymous):

since both sides are being divided by 12 you can multiply BOTH sides to rid of the fraction leaving only \[4x + 20 = 3x - 9\]

OpenStudy (anonymous):

Do i combine like terms

OpenStudy (anonymous):

yes :] you are very smart i think your just not trying ;P

OpenStudy (anonymous):

I got it thanks ;)

OpenStudy (anonymous):

ok good